OpenStudy (anonymous):

Please! Help!(: Find the antiderivative by integrating the function f(x)=csc(4x)cot(4x)

zepdrix (zepdrix):

\[\large \int\limits \csc(4x)\cot(4x)dx\] Hmmm, let's try a `U substitution` and see what happens. If we let \(\large u=\csc(4x)\), I think it will lead somewhere nice. We could try letting the other term be our u, just for the sake of trial and error, since that is the process you would probably take on this :) It would end up being the wrong sub, but it's the process that matters. Taking the derivative of our substitution gives us,\[\large du=-4\csc(4x)\cot(4x)dx\] If you're not familiar with that derivative, then the last step maybe have been a little confusing. Understand what I've done so far? :o

OpenStudy (anonymous):

kind of? where did that -4 come from?

OpenStudy (anonymous):

\[\int\limits_{}^{} \csc(4\theta)\cot(4\theta) d \theta \] That was the real function

zepdrix (zepdrix):

\[\large \left(\csc x\right)' \qquad=\qquad -\csc x \cot x\] The 4 came from the chain rule. :D

OpenStudy (anonymous):

I have to evaluate the integral.

zepdrix (zepdrix):

Yes, that's what we're working on. I just used x's instead of theta's.

zepdrix (zepdrix):

I guess a substitution like that is really not necessary.. Because if you know this derivative,\[\large \left(\csc x\right)' \qquad=\qquad -\csc x \cot x\] Then certainly you know this integral,\[\large \int\limits -\csc x \cot x dx\] Ok my mistake, all we really need to do is deal with the 4theta inside.

OpenStudy (anonymous):

oh okay. so howd you get from \[\csc(4x)\cot(4x)\] to \[(cscx)'=-\csc(x) \cot(x)\]

OpenStudy (anonymous):

I'm so confused. :( haha

zepdrix (zepdrix):

This is just a derivative rule that I was mentioning,\[(cscx)'=-\csc(x) \cot(x)\] If you haven't learned that yet, then maybe we should go about doing this the more tedious way ~ Converting cosecant and cotangent to sines and cosines and then do a u-substitution. :D

OpenStudy (anonymous):

could we convert then try the u-sub.?

zepdrix (zepdrix):

\[\large \int\limits \csc(4\theta)\cot(4\theta) d \theta \qquad=\qquad \int\limits \frac{1}{\sin(4\theta)}\left(\frac{\cos(4\theta)}{\sin(4\theta)}\right)d \theta\] So converting to sines and cosines gives us,\[\large \int\limits \frac{\cos(4\theta)d \theta}{\sin^2(4\theta)}\]Following along so far? :)

OpenStudy (anonymous):

yes(:

zepdrix (zepdrix):

Letting,\[\large u=\sin(4\theta)\] Taking the derivative of our substitution gives us,\[\large du=4\cos(4\theta)d \theta\]

zepdrix (zepdrix):

So with our last result right there, we want to divide each side by 4,\[\large \frac{1}{4}du=\cos(4\theta)d \theta\]

OpenStudy (anonymous):

okay.

OpenStudy (anonymous):

where do we go from there?

zepdrix (zepdrix):

\[\large \color{purple}{u=\sin(4\theta)}\]\[\large \color{green}{\frac{1}{4}du=\cos(4\theta)d \theta}\] We'll use this information to rewrite our integral in terms of u.\[\large \int\limits\limits \frac{\color{green}{\cos(4\theta)d \theta}}{\color{purple}{\sin(4\theta)}^2} \qquad\to\qquad \int\limits\limits \frac{\color{green}{\frac{1}{4}du}}{\color{purple}{u}^2} \]

OpenStudy (anonymous):

woah. okay haha.

zepdrix (zepdrix):

We'll rewrite our u using rules of exponents, \[\large \frac{1}{4}\int\limits u^{-2}du\] And from here we can apply the Power Rule for Integration to do the integration portion.

zepdrix (zepdrix):

What anti-derivative does that produce? :)

OpenStudy (anonymous):

-2u ^-2+1?

zepdrix (zepdrix):

Umm I'm not sure where the -2 is coming from, in front.

OpenStudy (anonymous):

the power?

zepdrix (zepdrix):

Remember, with integrals, you divide by the NEW EXPONENT. Not the -2, but the -1 that we end up with.

OpenStudy (anonymous):

2u^-1

zepdrix (zepdrix):

Ignoring the 1/4 fraction for a sec, here is what we want to be doing, \[\large \int\limits u^{-2} \qquad=\qquad \frac{u^{-2+1}}{-2+1} \qquad=\qquad -u^{-1}\]

OpenStudy (anonymous):

ohhhh. haha.

OpenStudy (anonymous):

okay I see where I went wrong. I was applying the power rule as a derivative. lol

zepdrix (zepdrix):

Ah :)

zepdrix (zepdrix):

So it looks like we end up with,\[\large \frac{1}{4}\int\limits\limits u^{-2}du \qquad=\qquad -\frac{1}{4}u^{-1}\]We'll want to do a little work to undo our substitution, and then write it in terms of csc or cot if we're able to.

OpenStudy (anonymous):

so. \[-\frac{ 1 }{ 4 }\sin(4\theta)^{-1}\]

zepdrix (zepdrix):

\[\large -\frac{1}{4}\left(\frac{1}{\sin(4\theta)}\right)+C\]Yup looks good! Remember what 1/sin gives us?

OpenStudy (anonymous):

csc?

zepdrix (zepdrix):

\[\large -\frac{1}{4}\csc(4\theta)+C\] Yay good job! \c:/

OpenStudy (anonymous):

Omg that was so hard! lol thank you so much!!!(:

zepdrix (zepdrix):

np

OpenStudy (anonymous):

do you know if you could help me with a similar prob/? @zepdrix

zepdrix (zepdrix):

Sure, post it in a new thread. This one is too long :U